Circular No. 6874

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J0421+560 AND CI CAMELOPARDALIS B. A. Harmon and G. J. Fishman, Marshall Space Flight Center,

NASA; and W. S. Paciesas, University of Alabama, Huntsville, report for the BATSE team: “Hard x-rays from CI Cam have not been detected with BATSE since Apr. 2–3. The duration of this transient event was unusual in both its rapid rise to peak flux (∼ 0.1 day) and decay (e-folding decay timescale of ∼ 0.56 day). Both of these timescales are much shorter than those seen from other galactic transients observed with BATSE over the past seven years. The peak of the hard x-ray flux occurred between Mar. 31.91–Apr. 1.04 UT at a level of 1160 ± 80, 950 ± 75, 650 ± 70, and 290 ± 75 mCrab in the bands 20–30, 30–40, 40–50, and 50–70 keV, respectively. BATSE detects the source only marginally in the band 70–100 keV. The initial flux estimate given (IAUC 6856) was underestimated due to the relatively soft spectrum of the source (power-law number index of −3.9 ± 0.2). No evolution of the spectral shape was detected. The flux history of XTE J0421+560 = CI Cam and other sources (updated daily) observed with BATSE by the Earth occultation technique can be accessed

from http://www.batse.msfc.nasa.gov/data/occult/fluxhistories/
eomon.html.”

GRB 980329 M. R. Metzger, California Institute of Technology, writes: “Near-

infrared K-band images of the area surrounding the variable radio source detected by Taylor et al. (GCN 40; cf. IAUC 6872), possibly associated with GRB 980329 (IAUC 6853, 6854), were obtained on Apr. 6.27 and 8.28 UT with the Keck-I 10-m telescope. The source reported by Larkin et al. (GCN 51, coincident with the galaxy reported by Djorgovski et al. on GCN 41) was detected on both nights, at K = 21.4 ± 0.2 and 21.9 ± 0.4, respectively. Comparison with the magnitudes reported by Larkin et al. indicates that the source has faded by ∼ 1 mag between Apr. 2.3 and 8.28. This suggests that part of the K flux may be infrared afterglow of GRB 980329; this is consistent with fading, detections, and upper limits reported previously (GCN 41, 46, 48; IAUC 6864, 6866, 6868). Further measurements are suggested to estimate the infrared brightness of any underlying host.”

1998 April 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green